[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Dec 06, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Dec 6 15:09:53 UTC 2017


Here is today's summary of NHC computer operations:
--- Dave

----------

TAFB:
1. OPC/WPC had problems with their surface map vgfs for yesterday's 12Z  
maps. Problem was resolved on their end.

2. Gecko has had some recent issues with NAWIPS loading spfs. Doug looked  
into the issue and found no system problems. TSB can assist with  
forecasters that need to migrate spfs to the new satellite directory  
structure. SPFs with GOES-16 may need to wait until the satellite has  
reached its new location next week.


HSU:
None
Other items:
None
NHC RFCs:
NCO RFC-3945 - Doug implemented a change on OPAH to allow CIRA and  
WeatherPredict access.

There is a scheduled upgrade for ETSS and PETSS on WCOSS today.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171206/6c1ea11c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list